A certain system can experience three different types of defects. Let A, i1,2,3) denote the event that the system has a defect of type i. Suppose that the following probabilities are true. p(A1)-0. 10 P(A2)-0.08 P(As)-0.06 (a) Given that the system has a type 1 defect, what is the probabilty that it has a type 2 defect? (Round your answer to four decimal places.) (b) Given that the system has a type 1 defect, what is the probability that it has all three types of defects? (Round your answer to four decimal places.) (c) Given that the systern has at least one type of defect, what is the probability that it has exactly one type of defect? (Round your answer to four decimal places.) (d) Given that the system has both of the first two types of defects, what is the probability that it does not have the third type of detect? (Round your answer to four decimal places.)
Answer #1

(a)

P(A_1 cup A_2 ) = P(A_1) + P(A_2) - P(A_1 cap A_2)

Rightarrow P(A_1 cap A_2 ) = P(A_1) + P(A_2) - P(A_1 cup A_2)

P(Al กั A-) = 0.10 0.08-0.12 = 0.06

Given type 1 defect, probability that it has a type 2 defect = P(A2 | A1)

= P(Ai n A2)/P(A) 0.06/0.10 0.6

(b)

Given type 1 defect, probability that it has all type 2 defect = P(A_1 cap A_2 cap A_3 | A_1)

P(A_1 cap A_2 cap A_3 cap A_1) / P(A_1 ) = P(A_1 cap A_2 cap A_3) / P(A_1 )

0.01/0.10 0.1

(c)

PlAi n As) = P(Ai) + P(As) _ P(A1 U As) = 0.1 + 0.06-0.13 0.03

PlA2 n As) = P(A2) + P(Аз) _ P(A2 U As) = 0.08 + 0.06-0.12 0.02

Probability that the system has at least one type of defect = P(A1 U A2 U A3)

P(A_1) + P(A_2) + P(A_3) - P(A_1 cup A_2) - P(A_1 cup A_3) - P(A_2 cup A_3) + P(A_1 cup A_2 cup A_3 )

= 0.1 + 0.08 + 0.06 - 0.06 - 0.03 - 0.02 + 0.01

= 0.14

By Venn diagram, Probability that it has exactly one type of defect = P(A_1 cup A_2 cup A_3 ) - P(A_1 cap A_2) - P(A_1 cap A_3) - P(A_2 cap A_3) + 2P(A_1 cap A_2 cap A_3 )

= 0.14 -  0.06 - 0.03 - 0.02 + 2 * 0.01 = 0.05

Given that the system has at least one type of defects, probability that it has exactly one type of defect

= P(exactly one type of defect | at least one type of defects)

= P(exactly one type of defect and at least one type of defects) / P(at least one type of defects)

= 0.05 / 0.14 = 0.3571

(d)

P(A_1 cap A_2 cap A_3^c) = P(A_1 cap A_2) - P(A_1 cap A_2 cap A_3) = 0.06 - 0.01 = 0.05

Given that the system has both of the first two types of defects, probability that it does not have third type of defect

= PLA: Ai n A2)-PLA1 n A2 A5)/P(A1 n A2) = 0.05/0.06 = 0.8333